Grade upto college level Physical Chemistry

Chromium metal crystallizes with a body centered cubic lattice. The length of the unit cell edge is found to be 287 pm. Calculate the atomic radius. What would be the density of chromium in g/cm3?

Shane Macguire
12 Years agoGrade upto college level

1 Answer

Deepak Patra
12 Years ago
Hello Student,
Please find the answer to your question
For bcc lattice, r = √3 *a/4 = √3/4 * 287 = 124.27 pm
Now Density = n *at. wt./V * Av. No.. = n * at. wt./a3 *Av. No.
N = 2 for bcc; a = 287 * 10-10 cm
∴ Density = 2 *51.99/(287 *10-10)3 * 6.023 *1023 = 7.30 g/ml

Thanks
Deepak patra
askIITians Faculty